OpenStudy (anonymous):

f(x)={x^2-1,x≤0 2x-2, x>1 find the domain and range of f(x) find f(0) and f(4)

OpenStudy (anonymous):

\[f(0)=0^2-1=-1\] \[f(4) =2\times 4 +2=6\]

OpenStudy (anonymous):

you use the top rule for 0 since \[0\leq 0\] and the bottom rule for 4 since \[4>1\]

OpenStudy (anonymous):

Thank you and what about the range and the domain